How many pounds of Raisind @ $2.20/lb must be mixed with 15 lbs of nuts @ $4.00/lb to restult in a trail mix worth $ 3.00/lb?
Answer by music974(43)

Ok well first we must make a table like this...
Amt. $ Total
Raisins r 2.20 2.20r
Nuts 15 4.00 60
Total 15+r 3.00 (15+r)3
Now we have to write an equation which would be 2.2r+60=(15+r)3 then we would distribute the three to get 2.2r+60=45+3r Then subtract both sides by 45 to get 2.2r+15=3r. Now subtract each side by 2.2r to get 15=.8r. Now multiply each side by 10 to get rid of the decimal thus getting 150=8r. Then divide both sides by 8 to get r=18.5. So you would need 18.5 pounds of raisins. Hope this helps!!